### v1

# Translators Asymptotic to Planes

Abstract.
We prove that a vertical plane is the only complete translator, properly immersed in $\mathbb{R}^{3}$ and having finite topology, whose ends are asymptotic to vertical planes.
1. Introduction
Translators arise naturally in the study of mean curvature flow as self-similar solutions that move by translation. Their geometric and analytic properties have been studied extensively, see for instance [1, Chapter 13], mainly because of their role in understanding type-II singularities of mean curvature flow [3, 4].
Our main theorem classifies complete properly immersed translators with finite topology whose ends are asymptotic to vertical planes, see Definition $2.$
Theorem $1.$
Let $\Sigma$ be a complete translator properly immersed in $\mathbb{R}^{3}$ with finite topology whose ends are asymptotic to vertical planes. Then $\Sigma$ is a vertical plane.
Our study was inspired by Khan’s work on complete embedded translators with finite total curvature [5], which, among other things, shows that such surfaces satisfy the hypotheses of Theorem $1.$ We therefore have the following corollary of Theorem $1.$
Corollary 1.2.
Every complete translator embedded in $\mathbb{R}^{3}$ with finite total curvature is a vertical plane.
Our proof of Theorem 1.1 relies principally on an identity expressing the total mean curvature of any smooth compact translator as the flux of $e_{3}^{\top}$ across its boundary, see Lemma 2.1.
2. Preliminaries
A translator $\Sigma\subset\mathbb{R}^{3}$ is a smooth surface whose mean curvature vector $H$ satisfies
$H=e_{3}^{\perp}.$
Equivalently, the family of surfaces $\Sigma+te_{3}$ for $t\in\mathbb{R}$ is a solution to the mean curvature flow.
In the following lemma, we express the total mean curvature of any smooth compact translator as the flux of $e_{3}^{\top}$ across its boundary.
Lemma 2.1.
For a smooth compact translator $\Sigma$ with boundary $\partial\Sigma$ we have
(1) $\int_{\Sigma}|H|^{2}=\int_{\partial\Sigma}\langle\eta,e_{3}^{\top}\rangle,$
where $\eta$ is the outward conormal to $\partial\Sigma$.
Proof.
Integrate the identity
$\operatorname{div}_{\Sigma}e_{3}^{\top}=-\operatorname{div}_{\Sigma}e_{3}^{\perp}=\langle H,e_{3}^{\perp}\rangle=|H|^{2}$
and apply the divergence theorem. ∎
Finally, we define what it means for an end to be asymptotic to a vertical plane.
Definition $2.$
Let $\Sigma_{i}$ be an end of a translator $\Sigma$, i.e. a component of $\Sigma\setminus B_{r}(0)$ for some sufficiently large radius $r$. We say that $\Sigma_{i}$ is asymptotic to a vertical plane if there exists a vertical plane $\Pi_{i}$ such that the following holds:
$\Sigma_{i}=\operatorname{graph}u_{i}$ for a function $u_{i}:\Pi_{i}\setminus\Omega_{i}\to\mathbb{R}$, where $\Omega_{i}\subset\Pi_{i}$ is open and bounded, and $\|u_{i}\|_{C^{1}(\Pi_{i}\setminus B_{R}(0))}\to 0$ as $R\to\infty$.
3. Proof of Theorem 1.1
To prove Theorem 1.1 we produce an exhaustion of $\Sigma$ by smooth, compact regions such that the boundary integral in (1) converges to zero. This implies that $H$ vanishes identically and hence $\Sigma$ is a vertical plane.
We first prove a variant of Theorem 1.1 which imposes a slightly different hypothesis on the ends.
Proposition 3.1.
Let $\Sigma$ be a complete translator properly immersed in $\mathbb{R}^{3}$. Suppose there exists a radius $r>0$ such that:
  • $\Sigma\setminus B_{r}(0)$ has finitely many components, denoted $\Sigma_{i}$;
  • each component $\Sigma_{i}=\operatorname{graph}u_{i}$ for a function $u_{i}:\Pi_{i}\setminus\Omega_{i}\to\mathbb{R}$, where $\Omega_{i}\subset\Pi_{i}$ is open and bounded, and
    (2) $\lim_{R\to\infty}\int_{\Pi_{i}\,\cap\,\partial B_{R}(p_{i})}|Du_{i}|^{2}=0$
    for some $p_{i}\in\Pi_{i}$.
Then $\Sigma$ is a vertical plane.
Proof.
For each $\Sigma_{i}$, up to an ambient isometry we may assume $\Pi_{i}=\{x_{1}=0\}$ and $p_{i}=0$. We introduce polar coordinates for $\Pi_{i}$, conflating $(0,r\cos\theta,r\sin\theta)$ with $(r,\theta)$. We may then express points in $\Sigma_{i}$ as $(u_{i}(r,\theta),r,\theta)$. For all sufficiently large $R$ we define map
$\gamma_{i,R}(\theta):=(u_{i}(R,\theta),R,\theta),$
giving rise to a smoothly embedded closed curve
$\Gamma_{i,R}:=\gamma_{i,R}(S^{1})\subset\Sigma_{i}.$
Since there are only finitely many components $\Sigma_{i}$ and $\Sigma$ is properly immersed, for each large $R$ the collection of curves $\Gamma_{i,R}$ bounds a compact region in $\Sigma$, which we denote by $\Sigma_{R}$. We will demonstrate that
$\lim_{R\to\infty}\int_{\Gamma_{R,i}}\langle\eta,e_{3}^{\top}\rangle=0$
for each $i$, where $\eta$ is the outward unit conormal to $\partial\Sigma_{R}$. The claim then follows, since Lemma 2.1 implies that
$\lim_{R\to\infty}\int_{\Sigma_{R}}|H|^{2}=0$
and hence $H$ vanishes everywhere, meaning $\Sigma$ can only be a vertical plane.
For the following computation let $i$ and $R$ be fixed and to ease notation let us write $\gamma$ instead of $\gamma_{i,R}$, $\Gamma$ instead of $\Gamma_{i,R}$, etc. The outward conormal to $\Gamma$ is
$\eta(R,\theta)=\frac{\gamma^{\prime}(\theta)}{|\gamma^{\prime}(\theta)|}\times\nu(R,\theta).$
where $\nu$ is the upward unit normal to $\Sigma$, given by
$\nu=\frac{e_{1}-Du}{\sqrt{1+|Du|^{2}}}.$
We therefore have
$\int_{\Gamma}\langle\eta,e_{3}^{\top}\rangle$ $=\int_{0}^{2\pi}\left\langle\frac{\gamma^{\prime}(\theta)}{|\gamma^{\prime}(\theta)|}\times\nu(R,\theta),e_{3}\right\rangle|\gamma^{\prime}(\theta)|\,d\theta$
$=\int_{0}^{2\pi}\langle\gamma^{\prime}(\theta)\times\nu(R,\theta),e_{3}\rangle\,d\theta.$
Evaluating the integrand at $(R,\theta)$ gives
$\langle\gamma^{\prime}(\theta)\times\nu,e_{3}\rangle$ $=\langle\gamma^{\prime}(\theta),e_{1}\rangle\langle\nu,e_{2}\rangle-\langle\gamma^{\prime}(\theta),e_{2}\rangle\langle\nu,e_{1}\rangle$
$=\frac{1}{\sqrt{1+|Du|^{2}}}(R\sin\theta|D_{2}u|^{2}-R\cos\theta D_{3}uD_{2}u+R\sin\theta)$
$=R\sin\theta+R\cdot O(|Du(R,\theta)|^{2}),$
and hence we have
$\int_{\Gamma}\langle\eta,e_{3}^{\top}\rangle$ $=\int_{0}^{2\pi}R\sin\theta\,d\theta+\int_{0}^{2\pi}R\cdot O(|Du(R,\theta)|^{2})\,d\theta$
$=\int_{0}^{2\pi}R\cdot O(|Du(R,\theta)|^{2})\,d\theta.$
The hypothesis (2), expressed in polar coordinates, ensures that the right hand side tends to zero as $R\to\infty$. This completes the proof. ∎
Our aim now is to prove that any end which is asymptotic to a vertical plane automatically satisfies (2). This will make use of the following interior gradient estimate, which is obtained by a standard application of Bernstein’s technique. [^1] Møller pointed out to us that the same estimate can be obtained by the different arguments contained in [2, Proposition 3.10] and a standard interpolation inequality.
Lemma 3.2.
Let $\Sigma$ be a translator of the form $\operatorname{graph}u$, where $u:\Pi\cap\overline{B_{R}(p)}\to\mathbb{R}$, $\Pi$ is a vertical plane containing the point $p$, and we assume $R\geq 1$. If $|Du|^{2}\leq\frac{1}{4}$ holds everywhere in $\Pi\cap\overline{B_{R}(p)}$ then we have
$|Du(p)|^{2}\leq\frac{400}{R}\max_{\Pi\,\cap\,\partial B_{R}(p)}u^{2}.$
Proof.
Up to an ambient isometry we may assume $\Pi=\{x^{1}=0\}$. A unit normal to $\Sigma$ is given by
$\nu:=\frac{e_{1}-Du}{\sqrt{1+|Du|^{2}}},$
and we have
$\langle H,\nu\rangle=\bigg(\delta^{ij}-\frac{D^{i}uD^{j}u}{1+|Du|^{2}}\bigg)\frac{D_{i}D_{j}u}{\sqrt{1+|Du|^{2}}}=:a^{ij}\frac{D_{i}D_{j}u}{\sqrt{1+|Du|^{2}}}$
and
$\langle e_{3},\nu\rangle=-\frac{D_{3}u}{\sqrt{1+|Du|^{2}}}.$
Therefore, the translator condition for $\Sigma$ is equivalent to
$a^{ij}D_{i}D_{j}u+D_{3}u=0.$
Here and throughout the proof indices range over $\{2,3\}$.
Let us write $L:=a^{ij}D_{i}D_{j}+D_{3}$. We then have
$L\frac{u^{2}}{2}=a^{ij}D_{i}uD_{j}u=\frac{|Du|^{2}}{1+|Du|^{2}}.$
and, using the Cauchy–Schwarz inequality,
$L\frac{|Du|^{2}}{2}$ $=-D^{k}uD_{k}a^{ij}D_{i}D_{j}u+a^{ij}D_{i}D^{k}uD_{j}D_{k}u\geq\bigg(1-5\frac{|Du|^{2}}{1+|Du|^{2}}\bigg)|D^{2}u|^{2}.$
Let $\varphi\geq 0$ be a cutoff function and $\Lambda>0$ a constant, both to be chosen later. From the above computations we see that the auxiliary quantity $G:=\varphi^{2}\frac{|Du|^{2}}{2}+\Lambda\frac{u^{2}}{2}$ satisfies
$LG$ $=\frac{|Du|^{2}}{2}L\varphi^{2}+a^{ij}D_{i}\varphi^{2}D_{j}|Du|^{2}+\varphi^{2}L\frac{|Du|^{2}}{2}+\Lambda L\frac{u^{2}}{2}$
$\geq\frac{|Du|^{2}}{2}L\varphi^{2}+a^{ij}D_{i}\varphi^{2}D_{j}|Du|^{2}$
$\qquad+\bigg(1-5\frac{|Du|^{2}}{1+|Du|^{2}}\bigg)\varphi^{2}|D^{2}u|^{2}+\frac{\Lambda|Du|^{2}}{1+|Du|^{2}}.$
Using $|a|\leq 2$ we estimate
$L\varphi^{2}\geq-4\varphi|D^{2}\varphi|-4|D\varphi|^{2}-2\varphi|D_{3}\varphi|$
and
$a^{ij}D_{i}\varphi^{2}D_{j}|Du|^{2}$ $\geq-8\varphi|D\varphi||Du||D^{2}u|\geq-\frac{1}{2}\varphi^{2}|D^{2}u|^{2}-32|D\varphi|^{2}|Du|^{2},$
leading to
$LG$ $\geq-|Du|^{2}(2\varphi|D^{2}\varphi|+34|D\varphi|^{2}+\varphi|D_{3}\varphi|)$
$\qquad+\bigg(\frac{1}{2}-5\frac{|Du|^{2}}{1+|Du|^{2}}\bigg)\varphi^{2}|D^{2}u|^{2}+\frac{\Lambda|Du|^{2}}{1+|Du|^{2}}.$
Let $\varphi$ be chosen so that $\varphi(p)=1$, $\varphi\equiv 0$ on $\partial B_{R}(p)$, $0\leq\varphi\leq 1$ and
$R|D\varphi|\leq 2,\qquad R^{2}|D^{2}\varphi|\leq 8.$
In addition, let $\Lambda=400R^{-1}$. Inserting these choices above and using $R\geq 1$, we finally arrive at
$LG$ $\geq\bigg(\frac{400}{1+|Du|^{2}}-200\Bigg)R^{-1}|Du|^{2}+\bigg(\frac{1}{2}-5\frac{|Du|^{2}}{1+|Du|^{2}}\bigg)\varphi^{2}|D^{2}u|^{2}.$
The right-hand side is nonnegative in $B_{R}(p)$ since we are assuming $|Du|^{2}\leq\frac{1}{4}$, so by the maximum principle
$\frac{|Du(p)|^{2}}{2}\leq\max_{B_{R}(p)}G=\max_{\partial B_{R}(p)}G=\frac{400}{R}\max_{\partial B_{R}(p)}\frac{u^{2}}{2}.$
We now complete the proof of our main theorem.
Proof of Theorem $1.$
Let $\Sigma\subset\mathbb{R}^{3}$ be a complete translator satisfying the hypotheses of Theorem $1.$ We proceed by verifying that each one of the finitely many annular ends satisfies equation (2), so that the claim follows from Proposition 3.1.
Consider an annular end $\Sigma_{i}=\operatorname{graph}u_{i}$ and fix a point $p_{i}\in\Pi_{i}$ for reference. Since, by hypothesis, we have
$\sup_{\Pi_{i}\setminus B_{R}(p_{i})}|Du_{i}|\to 0\qquad\text{as}\qquad R\to\infty,$
we certainly have $|Du_{i}|\leq 1/4$ in the region $\Pi_{i}\setminus B_{R/2}(p_{i})$ whenever $R$ is sufficiently large. Therefore, given any $q\in\Pi_{i}\setminus B_{R}(p_{i})$, by Lemma 3.2 we have
$|Du_{i}(q)|^{2}\leq\frac{800}{R}\max_{\partial B_{R/2}(q)}u_{i}^{2}$
for sufficently large $R$. Since $q\in\Pi_{i}\setminus B_{R}(p_{i})$ was chosen arbitrarily we conclude that
$\sup_{\Pi_{i}\setminus B_{R}(p_{i})}|Du_{i}|^{2}\leq\frac{800}{R}\sup_{\Pi_{i}\setminus B_{R/2}(p_{i})}u_{i}^{2}$
for sufficiently large $R$. But we also assume
$\sup_{\Pi_{i}\setminus B_{R/2}(p_{i})}u_{i}^{2}\to 0\qquad\text{as}\qquad R\to\infty,$
so this shows that
$\sup_{\Pi_{i}\setminus B_{R}(p_{i})}|Du_{i}|^{2}\leq o(R^{-1})$
and hence
$\int_{\Pi_{i}\,\cap\,\partial B_{R}(p_{i})}|Du_{i}|^{2}=\int_{0}^{2\pi}R|Du_{i}|^{2}\,d\theta=o(1)$
as $R\to\infty$, which is (2). ∎

### v2

@@
 Let $\Sigma$ be a complete translator properly immersed in $\mathbb{R}^{3}$ with finite topology whose ends are asymptotic to vertical planes. Then $\Sigma$ is a vertical plane.
+We note that since $\Sigma$ is not assumed to be embedded in Theorem 1.1, its asymptotic vertical planes need not be parallel.
 Our study was inspired by Khan’s work on complete embedded translators with finite total curvature [5], which, among other things, shows that such surfaces satisfy the hypotheses of Theorem $1.$ We therefore have the following corollary of Theorem $1.$
@@
 Our proof of Theorem 1.1 relies principally on an identity expressing the total mean curvature of any smooth compact translator as the flux of $e_{3}^{\top}$ across its boundary, see Lemma 2.1.
+Remark 1.3.
+After originally posting this article we learned of the work [6], by Neves and Tian, concerning the classification of Lagrangian translators.
+Some elements of our proof are similar to theirs—in particular, they derive the same flux formula for the total mean curvature and utilise it to prove flatness in their setting.
+Acknowledgements. We are grateful to Felix Schulze for bringing to our attention the reference [6].
 2. Preliminaries
@@
 We may then express points in $\Sigma_{i}$ as $(u_{i}(r,\theta),r,\theta)$.
-For all sufficiently large $R$ we define map
+For all sufficiently large $R$ we define a map
 $\gamma_{i,R}(\theta):=(u_{i}(R,\theta),R,\theta),$
@@
 We will demonstrate that
-$\lim_{R\to\infty}\int_{\Gamma_{R,i}}\langle\eta,e_{3}^{\top}\rangle=0$
+$\lim_{R\to\infty}\int_{\Gamma_{i,R}}\langle\eta,e_{3}^{\top}\rangle=0$
 for each $i$, where $\eta$ is the outward unit conormal to $\partial\Sigma_{R}$. The claim then follows, since Lemma 2.1 implies that
@@
 This will make use of the following interior gradient estimate, which is obtained by a standard application of Bernstein’s technique.
-[^1] Møller pointed out to us that the same estimate can be obtained by the different arguments contained in [2, Proposition 3.10] and a standard interpolation inequality.
+[^1] We note that Gama, Martín and Møller have independently obtained the same estimate by different arguments—it is implied by [2, Proposition 3.10] via a standard interpolation inequality.
 Lemma 3.2.